Gud afternoon.please be so kind to assist me with the following:
1.solve the following equation
1.1. 6log4n+3log4n-3=log4n+2log4n
1.2. (x-15)3/2=27
1.3.log(x+10)=2
2. A survey carried out in a certain village that 1/2 of the people who were interviewed are youth,1/3 are people living with disability,while 1/5 of the remainder are pensioners and the remaining 2500 villagers did not take part in the survey.
2.1. show all the necessary calculation to determine the total number of people living in that village.
2.2. determine the difference in number between the people living with disability and pensioners.

Answer by Boreal(15235)

1.2 3(x-15)=54; 3x-45=54;
1.3 log (x+10)=2; raise everything to the 10th power x+10=100, x=90
Regarding 2.
This to me is unclear
50% interviewed are youth
33% are living with disability
20% of the remainder (that would be 1/5 of 1/6 or 1/30) are pensioners, and the remaining 2500 villagers did not take part in the survey. Since I don't have any guide to the number of people who were interviewed, only the results, nor the relationship between the number interviewed and the number not interviewed, I can't go further.
disability and pensioners I can do the fraction and the numbers for both are above. To get the total number of people living in the village, I need the information mentioned above. One number for say the youth would give me the number surveyed.